# Removing Reflections from RAW Photos

### v2

@@ 5.3 Upsampling
We compare to V-DESIRR [36] in Tab. 2 by upsampling the ground truth (GT) and using the base model (E2E). For best E2E performance, we fine tuned our upsampler and V-DESIRR with the base model. Our method performs best on $t$ and $r$ (ours). Cycle consistency loss improves V-DESIRR (+C), so we used this for E2E. We ablated the upsampler masking operations by using only the finisher head (Ours-NM); performance degraded almost to match V-DESIRR.
Comparing on GT images, Fig. 6 and Fig. A17, V-DESIRR produces strong artifacts, even after fine tuning (adding cycle-consistency losses did not help).
Inference of our E2E upsampler, up to preview size $1024\times 1364$, takes $4.52\mathrm{s}$ and $6.53\mathrm{s}$ on our MacBook and iPhone (hardware details are in Sec. 5.2).
Discussion. V-DESIRR amplifies errors at low resolutions by repeatedly upsampling its previous output images. By contrast, our model masks and copies the mixture features $\phi_{M}$ to the outputs $(T,R)$. This direct copy reduces error propagation. Errors can still occur when features that are not present in the low resolution inputs become visible at the next level upward (see Fig. A18). In such cases the low resolution $(t,r)$ cannot guide upsampling of such features, and the upsampler must infer the high resolution image to which the features belong.
@@ 5.4 Reflection editing
In Fig. 7 and Fig. A19 we show that the predicted reflection facilitates aesthetic editing and error correction. In Fig. 7, the reflection color and spatial arrangement is modified. Error correction is shown in Fig. A19, Sec. F.4. Edits were made in Photoshop using the tone-mapped $t$ and $r$ images, and the “Linear Dodge (Add)” layer blend mode. We found that this is often sufficient, but editing tools should blend directly in linear RGB to enhance realism.
@@ 6 Conclusion
We have described a reflection removal system that is trained solely on mixture images $m$ that are created with a photometrically and geometrically accurate reflection simulation that uses scene-referred images. Moreover, we search among millions of these $m$ for well-exposed, visible, and visually interpretable $m$ that encode natural priors. Our system comprises a novel base model and upsampler that perform well on real images, even though we have not trained it on them.
Since Farid and Adelson [12], many signals have been used to remove reflections. We add illuminant color and contextual photos. Our system operates on RAW images end-to-end, which further and aids removal. Lastly, our system also has some ability to remove lens flares, though they are not in our dataset, Fig. A16. Flare removal systems might therefore be pre-trained to remove reflections and thus contend with the difficulty of capturing realistic lens flares.
Supplementary Material Eric Kee Adam Pikielny Kevin Blackburn-Matzen Marc Levoy
@@ Appendix A Photometric reflection synthesis
As part of our photometric reflection synthesis pipeline, Func. 1, we compute a new exposure and white balance for the simulated mixture image, $m$, using Func. A2 and Func. A3, respectively. These functions follow ACR color processing, and use methods in the Adobe DNG SDK, Appendix G. The ACR color processing that produces XYZ source images is specified in Func. A4 and discussed in Appendix G.
@@ A.1 White balancing
To compute a new white balance within Func. A3, we use the C5 white balancer of Afifi et al. [3]. C5 white balances an input image by using an additional $n$ sample images that were captured from the same camera. We therefore cache $n$ samples for each camera in the dataset of RAW images, and remove all images for which there were not $n=7$ samples from the camera ($7$ is the C5 default).
Function A2 Compute the exposure of a simulated mixture $m$.
Input: A simulated mixture image $m$ and its associated component images $(t,r)$
@@
1:  Compute the WhiteXY of $t$. {SDK Func. A10}
2:  Compute the transform XYZ_to_sRGB using WhiteXY. {SDK Func. A12}
@@
6:     Compute the target value $\tau=$ sRGB_to_linear_sRGB($0.4$). {SDK Func. A18}
@@
White balancing with C5 requires that simulated mixtures $m=t+r$ in XYZ be transformed into camera color space. In Func. A3 we use the XYZ_to_CAM transform associated with the RAW source image of $t$ to simulate a camera from which the mixture was captured, since $t$ typically dominates $r$ in the sum due to attenuation by the glass (see Appendix B). The white balancer produces a new white point WhiteCAM_awb in camera color space. We then follow ACR color processing of white points in camera color space by transforming WhiteCAM_awb into XY coordinates and computing a new XYZ_to_CAM transform using DNG SDK Func. A8. This new transform into XYZ is then composed with Bradford adaptation (Func. A3 line 5) to construct a single, linear white balancing transform that operates on images in XYZ space (line 6).
Function A3 Compute the white balance transform.
@@
1:  Compute the XYZ_to_CAM using the WhiteXY of $t$. {SDK Func. A8}
@@
4:  Compute XYZ_to_CAM_awb and WhiteXY_awb from WhiteCAM_awb. {Func. A11+A8 or A9}
5:  Compute XYZ_to_XYZ_D50 from WhiteXY_awb. {SDK Func. A14}
@@
Over a large scale dataset, white balancer failures inevitably occur. These are handled by culling $m$ if the new white point is extremely different from the as-shot WhiteXY of $t$. Extreme changes to the true white point are not common because reflections that are of practical interest are either transparent, localized, or both. The new XY white point (WhiteXY_awb) can be further restricted to lie on the Planckian locus, and we found this to be sufficient for our source images, which were captured under typical illuminants. Projection from WhiteCAM_awb to WhiteXY_awb can be done using ACR Func. A11 and A8, or Func. A9 with the Planckian constraint, as noted in Func. A3, line 4.
In summary, the white balance computed in Func. A3 is a linear transform, which we denote XYZ_to_XYZ_awb, that composes three ACR color transforms: 1) it maps images into the camera color space of $t$, 2) it maps back to XYZ under a new white point, and 3) it applies Bradford adaptation to the D50 illuminant. In the simulation (Func. 1) this transform is applied to $m$, $t$, $r$, and $c$ so they are interpreted with respect to the same white point, lines 8-9.
@@ Appendix B Geometric reflection synthesis
Figure A10: Simulated geometric properties at extreme values. Blurs are typically subtle.
In particular, we synthesize transmission images $t=T(i)$ and reflection images $r=R(j)$ as functions of $i$ and $j$ that appropriately model Fresnel attenuation, perspective projection, double reflection, and defocus blur. Examples are shown in Fig. A10. We omit from $T$ effects related to global color, dirt, and scratches since existing photo editing tools are well equipped to correct them after reflection removal.
@@ B.1 Fresnel attenuation
$\alpha$ $=\frac{1}{2}(\alpha_{\perp}+\alpha_{\parallel})$
$\alpha_{\perp}$ $=\frac{\sin^{2}(\theta_{\mathrm{i}}-\theta_{\mathrm{i}}^{\prime})}{\sin^{2}(\theta_{\mathrm{i}}+\theta_{\mathrm{i}}^{\prime})}$
$\alpha_{\parallel}$ $=\frac{\tan^{2}(\theta_{\mathrm{i}}-\theta_{\mathrm{i}}^{\prime})}{\tan^{2}(\theta_{\mathrm{i}}+\theta_{\mathrm{i}}^{\prime})}$
where $\theta_{\mathrm{i}}^{\prime}=\arcsin(\frac{1}{\kappa}\sin\theta_{\mathrm{i}})$, and $\kappa$ is the refractive index of glass. For $\theta_{\mathrm{i}}\in[0^{\circ}$, $45^{\circ}]$, Fresnel attenuation factors account for up to $-4$ stops (underexposure), and gradually strengthen to $-1$ stop at $83^{\circ}$.
To specify $\theta_{\mathrm{i}}$, we define images $r=R(j)$ as originating from a mirror surface, with incident rays reaching the camera by the law of reflection. In the next section we decribe how to simulate a diversity of practical geometric configurations of the glass and camera to construct $\theta_{\mathrm{i}}$ and thus compute $\alpha$. Glass also attenuates the transmission $t=T(i)$ by $1-\alpha$. This is typically close to $1$, but at extreme angles it creates a visible darkening effect.
@@ B.2 Camera projection
Inclination angle $\phi$. Most glass is approximately vertical, so if the viewpoint of $t$ looks upward, the viewpoint of $r$ should as well. We use a pose estimator [18], and augment the search for realistic pairings $(t,r)$, Sec. 3.1, by checking if $\phi_{t}-\phi_{r}=\Delta_{\phi}$ is below a maximum absolute value. In addition to this inclination discrepancy filter, images are culled if their inclination angle $\phi$ exceeds a threshold.
@@
Azimuth angle $\theta$. Most glass in consumer photography is roughly planar. We constrain the camera azimuth with respect to the glass so that the camera rays strike this plane (accounting for the FOV). We randomly sample $\theta\sim\mathcal{U}(-\theta_{\mathrm{max}},\theta_{\mathrm{max}})$.
@@ B.3 Defocus blur
Our physically based sampling procedure simulates reflections with a realistic amount of defocus blur for consumer photography. An example of a strongly blurred reflection is shown in Fig. A10. We however find that reflections are typically sharp, as Lei [27] also notes.
@@ B.4 HDR Environment sampling
The IBLs [14] are captured under a fixed white point, which allows for the color of the illuminant (i.e., its white balance) to be mixed correctly with the RAW data. We calibrate the exposure of these indoor IBLs by setting their median intensity to match the median value of all indoor RAW images (the median contends with saturated pixels). This cropped HDR image can be photometrically combined, and geometrically transformed using functions $T$ or $R$.
@@ B.5 Double reflection
Glass panes introduce multiple reflective surfaces that create a double reflection or “ghosting” effect. Shih et al. [40] ascribe this to the thickness of a single or double pane, and show shifts up to $4$ pixels for thicknesses in 3–10mm under some viewing distances, but double reflections are often much larger. Gaps between panes reach $20$mm as reported commercially, and each pane adds up to $7$mm. These multiple reflecting surfaces are also not necessarily parallel, uniformly thick, or flat as assumed in [40]. These factors produce significant double reflections even in modern windows, including when the camera is distant. We simulate these complex effects by adopting the geometric model of [40] and allowing a greater range of thicknesses, 8–20mm. We uniformly sample a glass thickness, physical viewing distance, and refractive index. These facilitate a ray tracing procedure, detailed below.
@@
Fig. A10 shows an example of a simulated double reflection, selected to show a case when the primary and secondary reflections are significantly shifted. We note that no doubling effect can occur along the direction of the glass surface normal because the rays that enter the glass re-emerge after internal reflection at the same location they entered. Thus double reflection fields that follow our geometric model (and the model of [40]), in which there are two perfectly parallel planes, must exhibit a radial pattern around the image of the glass surface normal. These patterns are not always apparent in practice, which suggests that the geometrical arrangement of glass surfaces that is described by Shih [40] omits important factors. We nonetheless adopt their model as being sufficient because visible reflections are typically localized to regions of an image, which obscures the presence or absence of a radial center.
@@ Appendix C The contextual photo
One arrangement of a primary and a contextual camera is shown in Fig. A11 (see caption for explanation). This specific arrangement of cameras is neither required nor typical in practice, but it reveals general geometric differences between the views of primary and contextual cameras. The view of the so-called reflection camera is translated by $2d$, twice the distance $d$ to the glass. Furthermore, if the contextual camera is rotated $180^{\circ}$ from the primary, the latter view will be in an opposite direction. At extreme rotations, the views will have little or no overlap.
Because the translation and rotation of the contextual camera view can differ significantly from the primary camera, it is difficult to simulate a contextual view using a dataset of image pairs $t=T(i)$ and $r=R(j)$ that are used to create mixture images $m$ from the perspective of a primary camera. In particular, content from $j$ should not be copied into a simulated contextual image $c$, as trained models could learn to cheat by searching for patches of $r$ that have the same perspective projection in $c$. Such patches will not be present in practice.
We create a scalably large dataset of contextual images $c$ by noting that $c$ will often contain no common content with $r$ unless the photographer is asked to
@@
Figure A11: The geometry of a primary and contextual camera view. In this figure the two views are co-located (black dot), and the latter is rotated 180 with respect to the former. Neither condition is required or typical. The contextual camera frustum is shown at the bottom ($\perp$ symbol); the primary camera frustum is not drawn. The reflection contains the scenery that would be captured by a virtual camera behind the glass (open circle, dashed $\perp$), equidistant to the glass wrt the primary camera, and swung azimuthally in the opposite direction. Note, the object (green circle) appears at the right edge of the contextual camera’s image (small green circle above it), but slightly left of center in the virtual camera view (small green circle near top of figure), and hence it is slightly right of center in the captured primary photo because the virtual camera is flipped left/right.
@@
point the camera at what they see in the reflection. We minimize such burden on the photographer, and define the contextual image $c$ as any image of the reflection scene that does not view the same parts of the scene as $r$. This definition allows contextual images to capture lighting information (sunlight, incandescent, etc.) and scene semantics (outdoor, indoor, city, nature, etc.) to aid reflection removal.
@@
To construct $c$, we crop reflection source images, $j$, into non-overlapping left/right or top/bottom squares $(j_{0},j_{1})$, and similarly for transmission source images $i$. This yields four pairs for simulation $(i_{a},j_{b}):(a,b)\in\{0,1\}^{2}$ for all $(i,j)\in\mathcal{D}$. The mixture and context images are $(m_{ab},c_{1-b})$, where $m_{ab}=\mathcal{C}(T(i_{a})+R(j_{b}))$, and $c_{1-b}=\mathcal{C}(j_{1-b})$. We fix the capture function $\mathcal{C}$ for both $m$ and $c$ to have the same white balance so $c$ can describe the color of the reflection scene in addition to its semantics.

### v3

@@ 5.3 Upsampling
We compare to V-DESIRR [35] in Tab. 3 by upsampling the ground truth (GT) and using the base model (E2E). [^6] Inference of our E2E upsampler, up to preview size $1024\times 1364$, takes $4.52\mathrm{s}$ and $6.53\mathrm{s}$ on our 2021 MacBook Prop and iPhone 14 Pro. For best E2E performance, we fine tuned our upsampler and V-DESIRR with the base model. Our method performs best (ours). Cycle consistency loss improves V-DESIRR (+C), so we used this for E2E. We ablated the upsampler masking operations by using only the finisher head (Ours-NM); performance degraded almost to match V-DESIRR.
Comparing on GT images, Fig. 5 and Fig. S11, V-DESIRR produces strong artifacts, even after fine tuning (adding cycle-consistency losses did not help).
Discussion. V-DESIRR amplifies errors at low resolutions by repeatedly upsampling its previous output images. Instead, our model masks and copies the high-res mixture features $\phi_{M}$ to the output $T,R$. This direct copy reduces error propagation. Errors can still occur when features that are not present in the low resolution inputs become visible at the next level upward (e.g. Fig. S14) because the low resolution $t,r$ cannot guide upsampling of such features, and the
Figure 6: Reflection editing.
upsampler must infer the high resolution image to which the features belong.
@@ 5.4 Reflection editing
In Fig. 6 and Fig. S15 we show that the predicted reflection facilitates aesthetic editing and error correction. In Fig. 6, the reflection color and spatial arrangement is modified. Error correction is shown in Fig. S15, Sec. F.4. Edits were made in Photoshop using the tone-mapped $t$ and $r$ images, and “Linear Dodge” blend mode (but linear blending would be ideal).
@@ 6 Conclusion
We have described a de-reflection system that is trained solely on images from a photometrically and geometrically accurate simulation. Moreover, we have imbued these images with natural priors by searching among millions of them for well-exposed and visually interpretable cases. This RAW simulation dramatically improves results, more than prior model variations, and enables our models to perform well on real images without training on them.
Since Farid and Adelson [12], many cues have been used for de-reflection. We add illuminant color and context photos, and use RAW images end-to-end. Our models are thus sensitive and can uncover hidden reflections, Fig. S9; privacy should be protected. Our system can also remove lens flares, though they are not in the dataset, Fig. S10. Flare removal systems might therefore be pre-trained to remove reflections, since it is difficult to capture real lens flares.
Supplementary Material
Contents
Supplemental sections
  1. Appendix A.
    Photometric reflection synthesis
    1. A.1.
      White balancing
  2. Appendix B.
    Geometric reflection synthesis
    1. B.1.
      Fresnel attenuation
    2. B.2.
      Camera projection
    3. B.3.
      Defocus blur
    4. B.4.
      HDR environment sampling
    5. B.5.
      Double reflection
  3. Appendix C.
    Contextual photos
  4. Appendix D.
    Data collection
    1. D.1.
      Mixture search
    2. D.2.
      Image collection
    3. D.3.
      Dataset settings
  5. Appendix E.
    Reflection removal methods
    1. E.1.
      Base model architecture
    2. E.2.
      Upsampler architecture
  6. Appendix F.
    Results
    1. F.1.
      Evaluation methodology
    2. F.2.
      Base model comparisons
    3. F.3.
      Upsampler comparisons
    4. F.4.
      Editing applications
  7. Appendix G.
    Adobe Camera RAW DNG SDK
Supplemental figures
  1. Fig. S1.
    Examples of simulated geometric properties
  2. Fig. S2.
    Overview of the reflection dataset
  3. Fig. S3.
    Contextual camera geometry
  4. Fig. S4.
    Base model architecture
  5. Fig. S5.
    Upsampler architecture
  6. Fig. S6.
    Result comparisons with ground truth
  7. Fig. S7.
    Base model results in-the-wild
  8. Fig. S8.
    Base model results in-the-wild (continued)
  9. Fig. S9.
    Results on subtle reflections
  10. Fig. S10.
    Results for lens flare removal
  11. Fig. S11.
    Upsampler model comparison
  12. Fig. S14.
    Upsampler model failures
  13. Fig. S15.
    Reflection editing for error recovery
Supplemental functions (simulation and DNG SDK)
  1. Func. A.1.
    compute_exposure()
  2. Func. A.1.
    white_balance()
  3. Func. G.
    extract_xyz_images()
  4. Func. G.
    cam_to_rgb()
  5. Func. G.
    stage3_black_level()
  6. Func. G.
    highlight_recovery()
  7. Func. G.
    find_xyz_to_cam()
  8. Func. G.
    xyz_to_cam_awb()
  9. Func. G - G.
    Additional ACR functions
@@ Appendix A Photometric reflection synthesis
As part of our photometric reflection synthesis pipeline, Func. 3.3, we compute a new exposure and white balance for the simulated mixture image, $m$, using Func. A.1 and Func. A.1, respectively. These functions follow ACR color processing, and use methods in the Adobe DNG SDK, Appendix G. The ACR color processing that produces XYZ source images is specified in Func. G and discussed in Appendix G.
@@ A.1 White balancing
To compute a new white balance within Func. A.1, we use the C5 white balancer of Afifi et al. [2]. C5 white balances an input image by using an additional $n$ sample images that were captured from the same camera. We therefore cache $n$ samples for each camera in the dataset of RAW images, and remove all images for which there were not $n=7$ samples from the camera ($7$ is the C5 default).
Function S1 Compute the exposure of a simulated mixture $m$.
Input: A simulated mixture $m$ and its associated component images $(t,r)$
@@
1:  Compute the WhiteXY of $t$. {SDK Func. G}
2:  Compute transform XYZ_to_sRGB using WhiteXY. {SDK Func. G}
@@
6:     Compute the target value $\tau=$ sRGB_to_linear_sRGB($0.4$). {SDK Func. G}
@@
White balancing with C5 requires that simulated mixtures $m=t+r$ in XYZ be transformed into camera color space. In Func. A.1 we use the XYZ_to_CAM transform associated with the RAW source image of $t$ to simulate a camera from which the mixture was captured, since $t$ typically dominates $r$ in the sum due to attenuation by the glass (see Appendix B). The white balancer produces a new
Function S2 Compute the white balance transform.
@@
1:  Compute the XYZ_to_CAM using the WhiteXY of $t$. {SDK Func. G}
@@
4:  Compute XYZ_to_CAM_awb and WhiteXY_awb from WhiteCAM_awb. {Func. G+G or G}
5:  Compute XYZ_to_XYZ_D50 from WhiteXY_awb. {SDK Func. G}
@@
white point WhiteCAM_awb in camera color space. We then follow ACR color processing of white points in camera color space by transforming WhiteCAM_awb into XY coordinates and computing a new XYZ_to_CAM transform using DNG SDK Func. G. This new transform into XYZ is then composed with Bradford adaptation (Func. A.1 line 5) to construct a single, linear white balancing transform that operates on images in XYZ space (line 6).
Over a large scale dataset, white balancer failures inevitably occur. These are handled by culling $m$ if the new white point is extremely different from the as-shot WhiteXY of $t$. Extreme changes to the true white point are not common because reflections that are of practical interest are either transparent, localized, or both. The new XY white point (WhiteXY_awb) can be further restricted to lie on the Planckian locus, and we found this to be sufficient for our source images, which were captured under typical illuminants. Projection from WhiteCAM_awb to WhiteXY_awb can be done using ACR Func. G and G, or Func. G with the Planckian constraint, as noted in Func. A.1, line 4.
In summary, the white balance computed in Func. A.1 is a linear transform, which we denote XYZ_to_XYZ_awb, that composes three ACR color transforms: 1) it maps images into the camera color space of $t$, 2) it maps back to XYZ under a new white point, and 3) it applies Bradford adaptation to the D50 illuminant. In the simulation (Func. 3.3) this transform is applied to $m$, $t$, $r$, and $c$ so they are interpreted with respect to the same white point, lines 8-9.
@@ Appendix B Geometric reflection synthesis
In particular, we synthesize transmission images $t=T(i)$ and reflection images $r=R(j)$ as functions of $i$ and $j$ that appropriately model Fresnel attenuation, perspective projection, double reflection, and defocus blur. We omit from $T$ effects related to global color, dirt, and scratches since existing photo editing tools are well equipped to correct them after reflection removal. Examples are shown in Fig. S1, and an overview of the synthetic images is shown in Fig. S2.
Figure S1: Simulated geometric properties at extreme values. Blurs are typically subtle.
@@ B.1 Fresnel attenuation
$\alpha$ $=\frac{1}{2}(\alpha_{\perp}+\alpha_{\parallel})$ (1)
$\alpha_{\perp}$ $=\frac{\sin^{2}(\theta_{\mathrm{i}}-\theta_{\mathrm{i}}^{\prime})}{\sin^{2}(\theta_{\mathrm{i}}+\theta_{\mathrm{i}}^{\prime})}$ (2)
$\alpha_{\parallel}$ $=\frac{\tan^{2}(\theta_{\mathrm{i}}-\theta_{\mathrm{i}}^{\prime})}{\tan^{2}(\theta_{\mathrm{i}}+\theta_{\mathrm{i}}^{\prime})}~{}~{},$ (3)
where $\theta_{\mathrm{i}}^{\prime}=\arcsin(\frac{1}{\kappa}\sin\theta_{\mathrm{i}})$, and $\kappa$ is the refractive index of glass. For $\theta_{\mathrm{i}}\in[0^{\circ}$, $45^{\circ}]$, Fresnel attenuation accounts for up to $-4$ stops (underexposure), and gradually strengthens to $-1$ stop for rays that glancingly strike the glass at $83^{\circ}$.
To specify $\theta_{\mathrm{i}}$, we define images $r=R(j)$ as originating from a mirror surface, with incident rays reaching the camera by the law of reflection. In the next section we decribe how to simulate a diversity of practical geometric configurations of the glass and camera to construct $\theta_{\mathrm{i}}$ and thus compute $\alpha$. Glass also attenuates the transmission $t=T(i)$ by $1-\alpha$. This is typically close to $1$, but at extreme angles it creates a visible darkening effect (Fig. S2, example $26$). Typical attenuation levels are shown in Fig. S2 in the column labeled “reflection.”
@@ B.2 Camera projection
Figure S2: Dataset overview ($2048$p). All images are simulations that use RAW sources. The numbers are referenced in the text.
Inclination angle $\phi$. Most glass is approximately vertical, so if the viewpoint of $t$ looks upward, the viewpoint of $r$ should as well. We use a pose estimator [18], and augment the search for realistic pairings $(t,r)$, Sec. 3.1, by checking if $\phi_{t}-\phi_{r}=\Delta_{\phi}$ is below a maximum absolute value. In addition to this inclination discrepancy filter, images are culled if their inclination angle $\phi$ exceeds a threshold. This aligns the horizons of $t$ and $r$, which introduces spatial priors, as illustrated in Fig. S2 examples $2$, $5$, $14$, $20$, and $24$.
@@
Azimuth angle $\theta$. Most glass in consumer photography is roughly planar. We constrain the camera azimuth with respect to the glass so that the camera rays strike this plane (accounting for the FOV). We randomly sample $\theta\sim\mathcal{U}(-\theta_{\mathrm{max}},\theta_{\mathrm{max}})$. The effect of this constraint can be most easily seen in Fig. S2 where highly oblique camera angles create spatially varying Fresnel attenuation across the reflection component (examples $6$, $7$, $14$, $19$, $23$, and $26$).
@@ B.3 Defocus blur
Our physically based sampling procedure simulates reflections with a realistic amount of defocus blur for consumer photography. An example of a strongly blurred reflection is shown in Fig. S1. We however find that reflections are typically sharp, as Lei [28] also notes. Typical defocus blurs are shown in Fig. S2; strong blurs are shown in examples $22$ and $23$.
@@ B.4 HDR Environment sampling
The IBLs [14] are captured under a fixed white point, which allows for the color of the illuminant (i.e., its white balance) to be mixed correctly with the RAW data. We calibrate the exposure of these indoor IBLs by setting their median intensity to match the median value of all indoor RAW images (the median contends with saturated pixels). This cropped HDR image can be photometrically combined, and geometrically transformed using functions $T$ or $R$. The effect of HDRs is shown in Fig. S2: reflected light sources, windows, etc. are produced by HDRs in examples $1$, $2$, $5$, $6$, $11$, $14$, $20$, $24$, and $26$.
@@ B.5 Double reflection
Glass panes introduce multiple reflective surfaces that create a double reflection or “ghosting” effect. Shih et al. [39] ascribe the effect of double reflection to the thickness of a single or double pane, and show shifts up to $4$ pixels for thicknesses in 3–10mm under some viewing distances, but double reflections are often much larger. Gaps between panes reach $20$mm as reported commercially, and each pane adds up to $7$mm. These multiple reflecting surfaces are also not necessarily parallel, uniformly thick, or flat as assumed in [39]. These factors produce significant double reflections even in modern windows, including when the camera is distant. We simulate these complex effects by adopting the geometric model of [39] and allowing a greater range of thicknesses, 8–20mm. We uniformly sample a glass thickness, physical viewing distance, and refractive index. These facilitate a ray tracing procedure, detailed below. Fig. S2 shows double reflections in the dataset; see examples $2$, $4$, $5$, $7$, $8$, $12$, $14$, and $15$.
@@
Fig. S1 shows an example of a simulated double reflection, selected to show a case when the primary and secondary reflections are significantly shifted. We note that no doubling effect can occur along the direction of the glass surface normal because the rays that enter the glass re-emerge after internal reflection at the same location they entered. Thus double reflection fields that follow our geometric model (and the model of [39]), in which there are two perfectly parallel planes, must exhibit a radial pattern around the image of the glass surface normal. These patterns are not always apparent in practice, which suggests that the geometrical arrangement of glass surfaces that is described by Shih [39] omits important factors. We nonetheless adopt their model as being sufficient because visible reflections are typically localized to regions of an image, which obscures the presence or absence of a radial center.
@@ Appendix C The contextual photo
One arrangement of a primary and a contextual camera is shown in Fig. S3 (see caption for explanation). This specific arrangement of cameras is neither required nor typical in practice, but it reveals general geometric differences between the views of primary and contextual cameras. The view of the so-called reflection camera is translated by $2d$, twice the distance $d$ to the glass. Furthermore, if the contextual camera is rotated $180^{\circ}$ from the primary, the latter view will be in an opposite direction. At extreme rotations, the views will have little or no overlap.
Because the translation and rotation of the contextual camera view can differ
@@
Figure S3: The geometry of a primary and contextual camera view. In this figure the two views are co-located (black dot), and the latter is rotated 180 with respect to the former. Neither condition is required or typical; they are shown to illustrate one possible geometric arrangement. The contextual camera frustum is shown at the bottom ($\perp$ symbol); the primary camera frustum is not drawn. The reflection contains the scenery that would be captured by a virtual camera behind the glass (open circle, dashed $\perp$), equidistant to the glass wrt the primary camera, and swung azimuthally in the opposite direction. Note, the object (green circle) appears at the right edge of the contextual camera’s image (small green circle above it), but slightly left of center in the virtual camera view (small green circle near top of figure), and hence it is slightly right of center in the captured primary photo because the virtual camera is flipped left/right.
significantly from the primary camera, it is difficult to simulate a contextual view using a dataset of image pairs $t=T(i)$ and $r=R(j)$ that are used to create mixture images $m$ from the perspective of a primary camera. In particular, content from $j$ should not be copied into a simulated contextual image $c$, as trained models could learn to cheat by searching for patches of $r$ that have the same perspective projection in $c$. Such patches will not be present in practice.
@@
We create a scalably large dataset of contextual images $c$ by noting that $c$ will often contain no common content with $r$ unless the photographer is asked to point the camera at what they see in the reflection. We minimize such burden on the photographer, and define the contextual image $c$ as any image of the reflection scene that does not view the same parts of the scene as $r$. This definition allows contextual images to capture lighting information (sunlight, incandescent, etc.) and scene semantics (outdoor, indoor, city, nature, etc.) to aid reflection removal.
@@
To construct $c$, we crop reflection source images, $j$, into non-overlapping left/right or top/bottom squares $(j_{0},j_{1})$, and similarly for transmission source images $i$. This yields four pairs for simulation $(i_{a},j_{b}):(a,b)\in\{0,1\}^{2}$ for all $(i,j)\in\mathcal{D}$. The mixture and context images are $(m_{ab},c_{1-b})$, where $m_{ab}=\mathcal{C}(T(i_{a})+R(j_{b}))$, and $c_{1-b}=\mathcal{C}(j_{1-b})$. We fix the capture function $\mathcal{C}$ for both $m$ and $c$ to have the same white balance so $c$ can describe the color of the reflection scene in addition to its semantics. Fig. S2 shows example contextual photos.